• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name nanda

    Log entry time 16:27:47 on March11,2005

    Entry number 141227


    Compton chicane beam tune was established this afternoon successfully. The beam quality is good. With 20uA beam at 4.6 GeV, photon signal and background rates (red) are 17 and 4 KHz respectively. Similar for electron (blue) rates. The photon spectrum looks reasonable (fig.2) .

    However, the bit in the DAQ hardware that tells the analysis code whether the laser polarization is left or right, is not being set correctly. Need access to troubleshoot this pattern unit.

    FIGURE 1

    FIGURE 2